OpenStudy (anonymous):

What is the slope of a line that passes through the point (−2, 3) and is parallel to a line that passes through (3, 7) and (−2, −8)? Numerical Answers Expected!

OpenStudy (anonymous):

line passing through (3,7) and (-2,-8) using two point formula take (x1,y1) = (3,7) (x2,y2) = (-2,-8) substitute in the equation solving it we get 3x-y = 18 given this line is parallel to the line passing through point (-2,3) as the line is parallel we can say it has same slope equation of the unknown line is 3x - y = c to find c substitute the point (-2,3) in the above eqn 3(-2) - 3 = c c = -9 therefore the eqn of the line is 3x-y = -9